# Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

### v1

@@ Abstract
Stratified Selective Sampling for Instruction Tuning
@@
Paramita Mirza1, Lucas Weber1, Fabian Küch1 1Fraunhofer IIS {first.last}@iis.fraunhofer.de
@@ 1 Introduction
Data selection has long been a central challenge in machine learning, aimed at curating datasets that maximise model performance. This has gained particular significance in the era of large language models (LLMs) (Albalak et al., 2024). LLM training typically spans three stages—pretraining, instruction-tuning, and alignment—with data selection playing a crucial role in each (e.g. Grattafiori et al., 2024b). However, the training objectives differ between each phase, and the goals of data selection also vary accordingly. While in pretraining the selection mechanisms are relatively coarse, relying on heuristics or lightweight classifiers to filter out the most undesirable data (Rae et al., 2021; Lee et al., 2022; Weber et al., 2024; Penedo et al., 2024), the post-training phase of instruction-tuning and subsequent alignment requires, in comparison, more rigorous selection strategies to guide the model toward desirable behaviors (Longpre et al., 2023; Conover et al., 2023; Wang et al., 2023; Grattafiori et al., 2024b).
Recent research on instruction tuning (IT) underscores the critical importance of data quality over sheer data quantity, establishing the benefits of IT data selection beyond merely reducing costs (see e.g. Zhou et al., 2023a; Liu et al., 2025; Qin et al., 2025), as larger datasets not only yield diminishing returns as they expand, but they can have detrimental effects on performance, especially with synthetic data (Ge et al., 2024; Diddee and Ippolito, 2025). On the other hand, given the abundance of available instruction data, one has to make a selection decision.
@@
Figure 1: An illustration of our three-stage pipeline: we first classify instructions into seven types, then score each sample for instruction difficulty ($f_{i}$) and response quality ($q_{i}$). We rank their combined scores ($p_{i}$) and sample top scoring examples within embedding-space clusters as well as overall, while maintaining fixed category proportions.
Manually curating effective instruction sets is both time-consuming and labour-intensive (Zhou et al., 2023a). A promising alternative is to automatically filter a small but diverse set of effective instructions from the extensive pool of available instruction data. Previous studies have proposed strategies to select “effective” instructions, focusing on three key dimensions of data assessment (Qin et al., 2025): i) difficulty/relevance, assessing whether a sample is trivial or challenging and how much it can therefore contribute to performance improvements, ii) quality, which refers to the usefulness and accuracy of responses, and iii) diversity, emphasising the scope of within-domain variability and cross-domain coverage of the data.
Prior work and its limitations.
Prior work highlights the value of complex or challenging examples for post-training (Li et al., 2024a; Liu et al., 2024a), but often relies on prompting LLMs to estimate complexity (Liu et al., 2024b; Lu et al., 2024) and lacks domain generalizability (Muennighoff et al., 2025). Similarly, quality scoring methods (e.g., Liu et al., 2024b; Ge et al., 2024) often ignore task-specific scoring needs (e.g., solutions’ correctness for math or coding problems), and some depend on large or proprietary models, raising cost and reproducibility issues (Chen et al., 2024). Domain-aware scoring strategies require a preprocessing step to bucket samples by task type. While some works mention this as part of data analysis (Ouyang et al., 2022a; Conover et al., 2023) or filtering pipelines (Grattafiori et al., 2024b), concrete methods and applications remain underexplored. Furthermore, as we detail in Section 2, only a few prior works consider all aspects of data assessment simultaneously (e.g., Liu et al., 2024b), and most evaluate on a single benchmark, limiting generalizability.
Our contributions.
In this paper, we propose a robust instruction-tuning data selection framework that simultaneously considers difficulty, quality and diversity, and we demonstrate its effectiveness across a wide range of benchmarks. Our key contributions are as follows: (i) a novel data sampling strategy for instruction-tuning that integrates task types (e.g., math, coding, generation), difficulty, and quality scores, while preserving both inter- and intra-class diversity through clustering and ranking; (ii) an efficient method for classifying IT data samples by task type; (iii) a new and efficient approach to estimate difficulty scores using model performance across diverse benchmarks; (iv) task-specific quality scoring strategies, including custom-designed scorers, particularly for constrained generation and coding tasks focusing on responses’ correctness; and (v) a large-scale evaluation spanning models of different families and sizes, with comparisons to strong baselines.
@@ 2 Related work
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024b), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality. On the other hand, model-inherent criteria form another group of approaches that rely on signals from the target model itself, including: LESS (Xia et al., 2024), which estimates data influence using gradient information; SelectIT (Liu et al., 2024a), which measures uncertainty via token probability, prompt variation and multiple models’ assessments; SHED (He et al., 2024), which clusters data and estimates impact using Shapley values; instruction-following difficulty (IFD, Li et al., 2024a) measuring discrepancies between the model’s intrinsic generation capability and its desired response; and the approach by Li et al. (2024b), which evaluates sample utility based on how much it reduces loss when used as an in-context example.
@@
Our approach combines both perspectives: we use domain-specific LLM-scorers to assess quality, and leverage various models’ performances across benchmarks to estimate difficulty—that is, how likely an average model is to fail on a given instruction. For diversity, our approach closely follows CaR (Ge et al., 2024), which clusters data points and samples one representative from each cluster.
IT data categorization.
@@
Previous work has proposed various task types (e.g., open QA, brainstorming, creative writing), to guide IT data collection from human annotators (Ouyang et al., 2022a; Conover et al., 2023). Grattafiori et al. (2024b) finetuned Llama 3 8B for coarse-grained (e.g., mathematical reasoning) and fine-grained (e.g., geometry and trigonometry) topic classification to help filter low-quality samples, though they provide little detail on the methodology or intended downstream use. Other efforts impose tags or domain taxonomy on IT data to ensure diversity (Lu et al., 2024; Muennighoff et al., 2025). Dong et al. (2024) study how data composition across tasks affects model performance, but assume that ShareGPT only contains general alignment tasks—we found that nearly 30% of it actually consists of coding tasks. To our knowledge, no prior work leverages IT data categorization to apply task-specific scoring strategies during selection.
@@ 3 Methods
Our goal is to determine a subset of an arbitrarily large-scale instruction dataset that is effective when finetuning a given pre-trained base model. We define an effective dataset as one which achieves high performance on a given set of general evaluation tasks, while requiring relatively few model parameter updates. Formally, let $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ with $d_{i}=(x_{i},y_{i})$ be the full IT dataset of size $N$, where $x_{i}$ represents an input sequence (i.e. an instruction) and $y_{i}$ represents the corresponding output (i.e. desirable model response). We wish to select a subset $\mathcal{D}^{\prime}\subset\mathcal{D}$ of size $m$ (with $m\ll N$) that is most effective for instruction tuning.
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024a; Liu et al., 2024a, b; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in a three-step pipeline as illustrated in Figure 1:
@@
    Scoring. Each input–output pair $(x_{i},y_{i})$ is scored using a category-specific quality scorer (yielding $f_{i}$) and a general-purpose difficulty scorer (yielding $q_{i}$); the results are combined into an overall preference score $p_{i}$.
@@ 3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022b), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023a) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
@@ 3.2 Scoring
Domain-agnostic difficulty scorer.
@@
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024b; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either often lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025) or are strongly influenced by spurious features (e.g. the widely used deita-complexity is strongly biased towards long sequences; see Figure 7 in Appendix A.2.1; Liu et al., 2024b).
@@
Our goal is to train a general and robust difficulty scorer for our data selection pipeline that predicts how likely it is for an average model to solve a data point incorrectly, independent of its category $l$.
@@
To source the training set $\mathcal{D}^{\text{diff}}$ for such a scorer, we collect 20k instruction-response pairs, evenly distributed across categories $l_{i}\in\mathcal{L}$ (data sources are listed in Table 4; proportions are shown in Figure 8). We evaluate every item with a heterogeneous pool of 18 instruction-tuned LLMs (see Table 5). Subsequently, we apply multiple preprocessing steps to the model scores: First, we normalise the item scores to the interval $[0,1]$ and remove items with a score of 0 across all models, as they are likely to contain noise or annotation errors. To mitigate potential skews in the model performance distribution, we convert the absolute scores into relative deviations with respect to the model’s mean performance on the corresponding category source dataset, by subtracting the average from the absolute score on each item. Ultimately, a difficulty target for every item is obtained by averaging over the model pool. We fine-tune a Qwen-3-8B backbone (Team, 2025) with a single-layer regression head to minimise the mean-squared error on this dataset (training details can be found in Appendix A.2.3).
Response quality scorer.
@@
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b). However, these strategies underestimate the diverse problem types within instruction tuning dataset, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category defined in Section 3.1, focusing on i) mathematical reasoning traces for Math ($q_{\mathrm{math}}$), ii) code snippets for Coding ($q_{\mathrm{code}}$), and iii) instruction-following capability for Generation and Brainstorming ($q_{\mathrm{if}}$). For Reasoning, Factual QA and Extraction, we employ the deita-quality scorer [^1] https://huggingface.co/hkust-nlp/deita-quality-scorer ($q_{\mathrm{deita}}$, Liu et al., 2024b), a finetuned LLaMA-13B for quality assessment,
@@
Process reward models (PRMs, Lightman et al., 2023; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We score Math data points using Qwen2.5-Math-PRM-7B (Zhang et al., 2025b). We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
@@
We design a quality scoring framework for Coding samples, drawing inspiration from Wadhwa et al. (2024). For each data point $(x_{i},y_{i})$, we leverage code-oriented LLMs to: (i) assess the functional correctness of the code snippet in $y_{i}$ with respect to the problem $x_{i}$, and (ii) produce a revised version that improves or fixes the original code (see Figure 13, Appendix A.3). The resulting score, $q_{\mathrm{code}}$, is based on the normalized Levenshtein similarity between lines of the original ($lo_{0},...,lo_{n}$) and revised ($lr_{0},...,lr_{m}$) code: $nls=(\max(n,m)-\mathrm{lev}(lo,lr))/\max(n,m)$, where $\mathrm{lev}(lo,lr)$ is the line-level Levenshtein distance. If the original code is functionally correct, we set $q_{\mathrm{code}}=nls$; otherwise, $q_{\mathrm{code}}=nls/2$. If no code snippet is present, we assign $q_{\mathrm{code}}=0.5$ if $y_{i}$ is judged correct, and 0.0 otherwise.
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses. [^2] https://huggingface.co/spaces/livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3 for details).
@@
Taking inspiration from the IFEval benchmark (Zhou et al., 2023b) which defines “verifiable constraints” such as length (“400 or more words”) and keyword (“without using the word sleep”) constraints, we design a response quality scorer based on the fraction of expressed constraints ($\mathcal{C}_{exp}$) adhered by the response ($\mathcal{C}_{true}$). First, we use an LLM annotator to identify $\mathcal{C}_{exp}$, which comprises (span, constraint type) pairs $\{(s_{i},c_{i})\}^{n_{exp}}_{i=1}$, with $s_{i}$ represents the textual span found and $c_{i}$ is the corresponding constraint label. For example, given the prompt “Write a funny blog post with 400 or more words about the benefits of sleeping in a hammock, without using the word sleep.”, $\mathcal{C}_{exp}=$(400 or more words, length), (without using the word “sleep”, keyword avoided), (funny blog post, writing type)}. A list of considered constraint types is shown in Figure 14 (Appendix A.4). Next, $\mathcal{C}_{exp}$ is passed to a constraint checker module, which performs two steps:
@@
This yields $\mathcal{C}_{true}=\{(s_{j},c_{j})\}^{n_{true}}_{j=1}$, with which we compute the quality score as $q_{\mathrm{if}}=n_{true}*(n_{true}/n_{exp})$, giving more incentives to responses adhering to more constraints. If $\mathcal{C}_{exp}$ is empty, we ask an LLM judge to evaluate whether the response i) addresses the user’s intent, while ii) respecting any constraints expressed in the prompt, and to provide a final $score$ (1–10), which we use to compute the score as $q_{\mathrm{if}}=score/10$.
Our analysis with the IFEval benchmark dataset containing sample responses from ten models as our test bed (see Table 8, Appendix A.4), shows that Qwen3-14B (Team, 2025) outperformed other medium-sized Instruct-LLMs as both LLM annotator and judge (see Table 9, Appendix A.4). It achieved a macro F1-score of 0.86 for identifying expressed constraints, a Pearson correlation coefficient of 0.523 at the instance-level, and 0.995 at the model-level, where it effectively replicated the IFEval model ranking.

### v2

# Laser: Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

@@ Abstract
Laser: Stratified Selective Sampling for Instruction Tuning
@@
Paramita Mirza1, Lucas Weber2, Fabian Küch2 1ScaDS.AI & TU Dresden, Germany, 2Fraunhofer IIS, Germany paramita.mirza@tu-dresden.de    {lucas.weber, fabian.kuech}@iis.fraunhofer.de
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Large language models (LLMs) can perform a wide range of text-based tasks through chat interfaces. Their generalist abilities stem from an extensive post-training phase, during which they are optimized to generate useful responses to user queries (Sanh et al., 2022; Mishra et al., 2022; Wei et al., 2021; Ouyang et al., 2022). The choice of training data in this phase has a major impact on model performance (e.g. Zhou et al., 2023a; Xia et al., 2024; Liu et al., 2024b; Chen et al., 2024; Ge et al., 2024).
Prior work identifies three key properties of effective instruction-tuning (IT) data (Zhou et al., 2023a): i) difficulty/relevance, reflecting how much a query contributes to learning (Li et al., 2024b; Liu et al., 2024a, c; Zhao et al., 2024b); ii) quality, the usefulness and accuracy of responses (Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c); and iii) diversity, the scope of within-domain variability and cross-domain coverage of the data (Ge et al., 2024; Lu et al., 2024). Although many approaches address one or more of these aspects, they often face limitations that make them difficult to apply in practice.
@@
Figure 1: Overview of our three-stage pipeline, Laser. Instructions are first classified into seven types, then routed and scored for difficulty ($f$) and quality ($q$). Their combined scores ($p$) are ranked, and top examples are sampled both within embedding clusters and overall, while preserving fixed category proportions.
Most existing methods cover only a subset of criteria (e.g., diversity in Chen et al. 2024; quality and diversity in Ge et al. 2024). To our knowledge, Deita (Liu et al., 2024b) is the only complete pipeline that integrates all three simultaneously. Beyond completeness, prior work can be further refined in terms of generality and robustness. Many heuristic-based methods cannot be applied to different tasks (Muennighoff et al., 2025) or domains (Zhou et al., 2023a). Broader approaches—that do not limit themselves explicitly to a single domain—are usually evaluated only on narrow domains or task-types (e.g., open-ended language generation), fine-tuned on a small set of models (typically from the Llama or Mistral family with 3B–8B parameters), or limited to a single source dataset. Outside these settings, they often prove brittle, failing to consistently outperform random sampling or simple heuristics such as a longest baseline (Diddee and Ippolito, 2025; Zhao et al., 2024a). Ultimately, many methods are expensive as they rely on large or closed-weight LLM judges (Chen et al., 2024; Lu et al., 2024).
We address these limitations with Laser (Label-Aware Scoring and clustERing), a robust, complete, and efficient data selection pipeline. [^1] Code, model access, and replication details will be released upon publication at https://github.com/Fraunhofer-IIS/LASER
Our key contributions are the following:
  1. (i)
    A novel, complete and robust data sampling strategy for instruction-tuning that integrates efficient scoring of difficulty and quality via routing while maintaining both inter- and intra-class diversity through stratified sampling and clustering;
  2. (ii)
    A new and general approach to estimate the difficulty of instructions;
  3. (iii)
    Task-specific quality scoring strategies, including custom-designed scorers, particularly for constrained generation and coding tasks focusing on responses’ correctness; and
  4. (iv)
    A large-scale evaluation spanning diverse model families and sizes, on a wide range of benchmarks covering various task types, with comparisons to strong baselines.
@@ 2 Related work
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024c), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as Deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality.
On the other hand, model-inherent criteria form another group of approaches that rely on signals from the target model itself, including: LESS (Xia et al., 2024), which estimates data influence using gradient information; SelectIT (Liu et al., 2024a), which measures uncertainty via token probability, prompt variation and multiple models’ assessments; SHED (He et al., 2024), which clusters data and estimates impact using Shapley values; instruction-following difficulty (IFD; Li et al., 2024b), measuring discrepancies between the model’s intrinsic generation capability and its desired response; and the approach by Li et al. (2024c), which evaluates sample utility based on how much it reduces loss when used as an in-context example.
@@
Our approach embraces external and model-inherent perspectives simultaneously: we use domain-specific LLM-scorers to assess quality, and leverage various models’ performances across benchmarks to estimate difficulty—that is, how likely an average model is to fail on a given instruction. For diversity, our approach closely follows CaR (Ge et al., 2024), which clusters data points and samples one representative from each cluster.
IT data categorisation.
@@
Previous work has proposed various task types (e.g., open QA, brainstorming, creative writing), to guide IT data collection from human annotators (Ouyang et al., 2022; Conover et al., 2023). Grattafiori et al. (2024) finetuned Llama 3 8B for coarse-grained (e.g., mathematical reasoning) and fine-grained (e.g., geometry and trigonometry) topic classification to help filter low-quality samples, though they provide little detail on the methodology or intended downstream use. Other efforts impose tags or domain taxonomy on IT data to ensure diversity (Lu et al., 2024; Muennighoff et al., 2025). Dong et al. (2024) study how data composition across tasks affects model performance, but assume that ShareGPT only contains general alignment tasks, while our analysis shows that nearly 30% of it actually consists of coding tasks. To our knowledge, no prior work leverages IT data categorisation to apply task-specific scoring strategies during selection.
@@ 3 Methods
Our goal is to determine a subset of an arbitrarily large instruction source dataset that is effective when finetuning a given pre-trained base model. We define an effective dataset as one which achieves high performance on a large and general set of evaluation tasks, while requiring relatively few model parameter updates. Formally, let $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ with $d_{i}=(x_{i},y_{i})$ be the full source dataset of size $N$, where $x_{i}$ represents an input sequence (i.e. an instruction) and $y_{i}$ represents the corresponding output (i.e. the desirable model response). We wish to select a subset $\mathcal{D}^{\prime}\subset\mathcal{D}$ of size $m$ (with $m\ll N$) that is most effective for instruction tuning.
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024b; Liu et al., 2024a, c; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in the three-step Laser pipeline as illustrated in Figure 1:
@@
    Scoring. Each input–output pair $(x_{i},y_{i})$ is scored using a general-purpose difficulty scorer (yielding $f_{i}$) and a category-specific quality scorer (yielding $q_{i}$); the results are combined into an overall preference score $p_{i}$.
@@ 3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023a) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
@@ 3.2 Scoring
3.2.1 Difficulty scorer
@@
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024c; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025, or the Deita complexity scorer as shown in Figure 10(b)) or are strongly influenced by spurious features (e.g. the widely used Deita-complexity is strongly biased towards long sequences; see Figure 11; Liu et al., 2024c).
@@
Our goal is to train a general and robust difficulty scorer that predicts how likely it is for an average model to solve a data point incorrectly, independent of its category $l$. We denote this difficulty score as $f$.
@@
To source the training set $\mathcal{D}_{\text{diff}}$ for such a scorer, we collect 20k instruction-response pairs, approximately equally distributed across categories $l_{i}\in\mathcal{L}$. We list the data sources in Table 13 and show the category proportions in Figure 8. Then, we evaluate every item with a heterogeneous pool of 18 instruction‑tuned LLMs (see Table 4). We continue by applying multiple preprocessing steps to the model scores: First, we normalise the item scores to the interval $[0,1]$ and remove items with a score of 0 across all models, as they are likely to contain noise or annotation errors and will not yield any valuable learning signal. We then go on to convert the absolute scores into relative deviations with respect to the model’s mean performance on the corresponding dataset, by subtracting the average from the absolute score on each item (illustrated in Figure 9). This step has the effect of weighing wrong responses of strong models more than wrong responses of weaker models (and vice versa). Further, it mitigates potential skews in the model performance distribution (e.g. if many models perform weakly on a specific dataset). Ultimately, we obtain difficulty targets $f$ by averaging over the model pool for every item.
We fine‑tune a Qwen‑3‑8B backbone (Team, 2025) with a single‑layer regression head to minimise the mean‑squared error on this dataset (training details can be found in Appendix A.2.2).
Difficulty scorer evaluation.
We evaluate the effectiveness and the validity of the difficulty scoring. For effectiveness, we sample 25k data points from $\mathcal{D}$ (as it is described in Section 4.1) either at random or following the difficulty scores. We then finetune Mistral-7B-v0.3 (Jiang et al., 2023) on the resulting datasets. Appendix A.2.3 shows how the difficulty scorer helps improve performance on benchmarks independently of their domain. For validity (i.e. whether we are really measuring some notion of difficulty), we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) and find high correlations with its human difficulty scores (details in Appendix A.2.4).
3.2.2 Quality scorer
@@
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c). However, these strategies underestimate the diverse problem types within instruction tuning data, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category defined in Section 3.1.
Deita quality scorer.
For Reasoning, Factual QA and Extraction, we employ the Deita quality scorer [^2] https://huggingface.co/hkust-nlp/deita-quality-scorer ($q_{\mathrm{deita}}$, Liu et al., 2024c), a finetuned LLaMA-13B for quality assessment.
@@
For Math data, we assess the soundness of mathematical reasoning using a specialised process reward model (Qwen2.5-Math-PRM-7B; Zhang et al., 2025b). Process reward models (PRMs; Lightman et al., 2023; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
@@
We design a quality scoring framework for Coding samples by drawing inspiration from Wadhwa et al. (2024). For each data point $(x_{i},y_{i})$, we leverage code-oriented LLMs to: (i) assess the functional correctness of the code snippet in $y_{i}$ with respect to the problem $x_{i}$, and (ii) produce a revised version that improves or fixes the original code (see Figure 20, Appendix A.3.2). The resulting score, $q_{\mathrm{code}}$, is based on the normalized Levenshtein similarity between lines of the original ($lo_{0},...,lo_{n}$) and revised ($lr_{0},...,lr_{m}$) code: $nls=(\max(n,m)-\mathrm{lev}(lo,lr))/\max(n,m)$, where $\mathrm{lev}(lo,lr)$ is the line-level Levenshtein distance. If the original code is functionally correct, we set $q_{\mathrm{code}}=nls$; otherwise, $q_{\mathrm{code}}=nls/2$. If no code snippet is present, we assign $q_{\mathrm{code}}=0.5$ if $y_{i}$ is judged correct, and 0.0 otherwise.
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses. [^3] https://huggingface.co/spaces/livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3.2 for details).
@@
For Generation and Brainstorming data, we implement an if-quality scorer. Influenced by the IFEval benchmark (Zhou et al., 2023b) which defines “verifiable constraints” such as length (“400 or more words”) and keyword (“without using the word sleep”) constraints, we design a response quality scorer based on the fraction of expressed constraints ($\mathcal{C}_{exp}$) adhered by the response ($\mathcal{C}_{true}$). First, we use an LLM annotator to identify $\mathcal{C}_{exp}$, which comprises (span, constraint type) pairs $\{(s_{i},c_{i})\}^{n_{exp}}_{i=1}$, with $s_{i}$ represents the textual span found and $c_{i}$ is the corresponding constraint label. For example, given the prompt “Write a funny blog post with 400 or more words about the benefits of sleeping in a hammock, without using the word sleep.”, $\mathcal{C}_{exp}=$ {(400 or more words, length), (without using the word “sleep”, keyword avoided), (funny blog post, writing type)}. A list of considered constraint types is shown in Figure 21 (Appendix A.3.1). Next, $\mathcal{C}_{exp}$ is passed to a constraint checker module, which performs two steps:
@@
This yields $\mathcal{C}_{true}=\{(s_{j},c_{j})\}^{n_{true}}_{j=1}$, with which we compute the quality score as $q_{\mathrm{if}}=n_{true}*(n_{true}/n_{exp})$, giving more incentives to responses adhering to more constraints. If $\mathcal{C}_{exp}$ is empty, we ask an LLM judge to evaluate whether the response i) addresses the user’s intent, while ii) respecting any constraints expressed in the prompt, and to provide a final scorescore (1–10), which we use to compute the score as $q_{\mathrm{if}}=score/10$.
Our analysis with the IFEval benchmark dataset containing sample responses from ten models as our test bed (see Table 6, Appendix A.3.1), shows that Qwen3-14B (Team, 2025) outperformed other medium-sized Instruct-LLMs as both LLM annotator and judge (see Table 7, Appendix A.3.1). It achieved a macro F1-score of 0.86 for identifying expressed constraints, a Pearson correlation coefficient of 0.523 at the instance-level, and 0.995 at the model-level, where it effectively replicated the IFEval model ranking.
Quality scorer evaluation.
Similar to the difficulty scorer, we demonstrate the effectiveness of our task-specific quality scorers for ranking-based data selection, particularly when compared to the general-purpose Deita quality scorer and random sampling. The corresponding results are presented in Appendix A.3.3.
Dataset #samples (#turns) HuggingFaceH4/ifeval-like-data 5K vicgalle/alpaca-gpt4 (Taori et al., 2023) 52K nvidia/OpenMathInstruct-2 (w/o augmented problems) (Toshniwal et al., 2024) 52K ai2-adapt-dev/flan_v2_converted (Longpre et al., 2023) 90K openbmb/UltraInteract_sft (Coding) (Yuan et al., 2024) 115K WizardLMTeam/WizardLM_evol_instruct_V2_196K (Xu et al., 2024) 143K theblackcat102/sharegpt-english 50K (392K) microsoft/orca-agentinstruct-1M-v1 (random subset) (Mitra et al., 2024) 200K (903K) all 707K (1.75M)
Table 2: Source dataset $\mathcal{D}$
Figure 2: Category distribution in IT datasets with varied sampling.
3.3 Sampling